OpenStudy (anonymous):

Solve the system 3x - 5y = 8 and 4x - 7y = 10.

OpenStudy (anonymous):

solve one of the equations for x and then plug that into the other equation

OpenStudy (anonymous):

x = (8/3) + (5/3)y plug that into 4((8/3) + (5/3)y) - 7y =10

OpenStudy (ragingsquirrel):

yeah... X=6 Y=2

hero (hero):

3x - 5y = 8 4x-7y = 10 3x - 8 = 5y 4x - 10 = 7y y = (3x-8)/5 y = (4x-10)/7 y = y 7(3x-8) = 5(4x-10) 21x - 56 = 20x - 50 21x - 20 = 56 - 50 x = 6 y = 2

OpenStudy (anonymous):

Thanks, but I think you're supposed to write it in the form of exponents.

OpenStudy (anonymous):

Oh, nevermind. I figured it out. Thanks!

hero (hero):

You mean x,y = (6,2)